OpenStudy (anonymous):

The numbers 0 through 9 are written on index cards. Three of the cards are randomly selected to form a 3-digit code. How many different codes are possible? Is it 388?

OpenStudy (anonymous):

On the 1st card, there are 9 possible digits you could pick up On the 2nd card, there are 8 possible digits you could pick up On the 3rd card, there are 7 possible digits you could pick up Total=9*8*7=504

OpenStudy (anonymous):

That's inaccurate.

OpenStudy (anonymous):

Please, re-check that.

OpenStudy (anonymous):

wouldn't it be 720

OpenStudy (anonymous):

Sorry, 10*9*8- I missed out the 0, so there are 10 possible digits. 720 overall

OpenStudy (anonymous):

Ah, thank you! Do you think you could care to do another one? :)

OpenStudy (anonymous):

Find the number of favorable outcomes in the sample space for the experiment. A number cube is rolled twice. The sum of the rolls is 7

OpenStudy (anonymous):

Would it be 12?

OpenStudy (anonymous):

..

OpenStudy (anonymous):

1 + 6, 6 + 1 , 5+2, 2+5 = 4 outcomes

OpenStudy (anonymous):

oops sorry I forgot 3 + 4, 4 + 3

OpenStudy (anonymous):

so there is 6 outcomes